OpenStudy (anonymous):

THE SIDE OF A PARALLELOGRAM IS 1/4 THE LENGTH OF ITS BASE. THE PARALLELOGRAM'S PERIMETER IS 125. WHAT IS THE BASE AND THE SIDE OF THE PARALLELOGRAM?

OpenStudy (anonymous):

Again, if you let the base of the parallelogram be b, and the side s, you're told,\[s=\frac{1}{4}b\]You're also told that the perimeter is 125; that is,\[2s+2b=125\]You can sub. the first equation for s into the second to find for b,\[2(\frac{1}{4}b)+2b=125 \rightarrow \frac{1}{2}b+2b=\frac{5}{2}b=125\]Multiply both sides by 2, then divide both sides by 5. You get,\[b=50\]Since \[s=\frac{1}{4}b \rightarrow s=\frac{50}{4}=\frac{25}{2}=12.5\]